OpenStudy (anonymous):

Factor using the GCF: 16m + 4n - 12p

OpenStudy (anonymous):

Right away 4 stands out to me :-)

OpenStudy (anonymous):

16 / 4 = ? , 4 / 4 = ? , 12 / 4 = ?